OpenStudy (briellshaw):

Solve the equation for x. HINT: What do you have to do with the denominators? x + \frac{3}{8} = \frac{17}{20} A. \frac{14}{12} B. \frac{19}{40} C. \frac{32}{20}

OpenStudy (anonymous):

Get the same denominator. \[\frac{3}{8}=\frac{15}{40}\] \[\frac{17}{20}=\frac{34}{40}\] Try and solve x now

OpenStudy (briellshaw):

is my answer correct is it c

OpenStudy (briellshaw):

@to

OpenStudy (anonymous):

It is not c @BriellShaw

OpenStudy (anonymous):

Try and set up the equation

OpenStudy (briellshaw):

so its b

OpenStudy (anonymous):

Correct :)